A van travels 1500 km south with an initial velocity of 90 km/hr. Find the distance in km/hr if its acceleration is 45 km/hr.what is its final velocity in m/s?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

A van travels 1500 km south with an initial velocity of 90 km/hr. Find the distance in km/hr if its acceleration is 45 km/hr.
-----
The distance is 1500 km, can't be in km/hr.
If you meant the speed at 1500 km:
---
s = at^2/2 + vot
1500 = 45t^2 + 90t
3t^2 + 6t - 100 = 0
t =~ 4.859 hours
v at 1500 km = 90 + 45t =~ 308.676 km/hr
-----------------------
what is its final velocity in m/s?
---
Convert 308.676 km/hr to m/sec.
------------------------
I'm not sure that's what you meant to ask.

Answer by ikleyn(52788)

.
Two of five unit names in this post are wrong.

The post, the problem and questions make no sense.